# Statistics for Differential Topological Properties between Data Sets with an Application to Reservoir Computers

### v1

@@ I Introduction
What does all this have to do with reservoir computing systems? The main link there is provided by Takens embedding theorem. We will show a simplified explanation of how attractors on manifolds in dynamical systems driving a reservoir computer can (under the right circumstances) provide an embedding of the drive system in the reservoir system. The tests for this possibility using only data will involve the statistics we mentioned above. To get to this point we start by introducing reservoir systems and clarifying which principles are important for understanding their dynamics as well as which explanations of their behavior are misleading.
@@ II Reservoir Computing
Reservoir computing was introduced in the early 2000’s in a few papers Jaeger (2001); Maass, Natschlaeger, and Markram (2002); Jaeger (2007); H. Jaeger (2004), which influenced the machine learning field more than the nonlinear dynamics field, at least at first. To put it simply, it was initially shown that the reservoir computer (RC) could be driven by an input of one component of a dynamical system (the drive) and it could be trained to reproduce the other (unseen by the reservoir) drive components. And, once trained it could do this accurately for signals from other initial conditions of the same drive system.
@@ III Basic structure of Reservoir Computer and Test Systems
We note that this process does not necessarily work for any input from the drive. Which ones are successful varies from drive to drive. The statistics we develop here can be useful in determining which combinations of drive signals work best. The outputs of the RC are signals from the nodes of the RC that are from the dynamical variables of the nodes. We extract all nodes’ signals, but it’s still somewhat an open question of which nodes of a subset might work just as well.
@@ IV Dynamics and Embeddings in a drive-
where ${\bf r}\in{\mathbbm{R}}^{NM}$. We note that using function composition iteratively in Eq.(2) and iterating on ${\bf r}(n)$ we can write Eq.(2) in an insightful way Jüngling . That is, we write,
@@
From the above Eq.(3) we see that the "next" time step ($n+1$) depends iteratively on the previous $n$ steps, assuming $n=0$ is the initial time and ${\bf{r}}(0)$ and ${\bf{x}}(0)$ the initial conditions. Eq.(3) is reminiscent of the form of Taken’s theorem Takens (1981) with respect to a dependence of the dynamics reconstruction on multiple time shifted terms. We can show that with a few more general assumptions Taken’s theory indeed does apply here.
In the case of continuous time evolution of the RC and drive we can get the same type of result as for Eq.(3), although it is a little more involved. Let $\psi_{t}$ be the flow of the drive system and $\Psi_{t}$ be the flow of the RC. The dynamics of the system is in the form of a skew product flow Kloeden and Yang (2021). That is the drive $\psi_{t}$ is autonomous, but the RC $\Psi_{t}$ is nonautonomous. The important property we will need is the cocycle property Kloeden and Yang (2021) of the flow on a nonautonomous system (the RC): $\forall\>s,t\in\mathbbm{R}^{+}\text{, }\bf{x}$ in the drive space, and $\bf{r}$ in the RC space, the following describes the flow:
@@
Then, iterating the above drive-RC embedding equations analogous to Eq.(3) can be written as,
@@
We note that the above explanation does not amount to a strict mathematical proof of the connection between the RC trajectory and some diffeomorphism of the drive attractor. We simply want to show that such a relation is reasonable and even likely. For more rigorous proofs of the RC-drive attractor relationship see the papers Allen Hart (2020); Grigoryeva, Hart, and Ortega (2021). We note that proofs for the theorems in these papers are at times limited to drive and RC systems with special properties (e.g. the necessity of contracting maps as a property of the dynamics). We emphasize that the above statements are not criticisms of the referenced papers. The proofs are involved and highly non-trivial. But so far the research on RCs suggests that the actual requirements for an embedding are less restrictive.
@@
In the next few sections we take the view that many drive-RC systems can generate time series data that can be used to test for the likelihood that the system has certain mathematical properties. This requires the development of statistics that we mentioned in the introduction.
@@ V Continuity Statistic
In this section on continuity and in the next section on differentiability, for simplicity of explanation we will treat the process of finding spatially nearby points as one of a simple search for the closest data points. This process is actually more subtle for dynamical time series. We will address the proper way to find nearby, but time-uncorrelated points below.
Searching for closest points uses a metric on both spaces which we assume to be the usual square-root of sum of squares of component differences. This automatically translates into a set of open sets on each space as defined in the usual way by the metric and, hence, to a metric-induced topology on both drive and RC data spaces. As a result we use the same metric to define the properties of continuity, differentiability, etc. without resorting to a topological definition in terms of function inverses, etc. The latter would add a more abstract approach, but it would be unnecessarily so.
@@
Note that this type of mapping preserves a lot of geometric properties between the manifolds. Many properties of ${\mathcal{M}}$ are also true of ${\mathcal{N}}$ and vice versa. Especially important is the fact that many dynamical properties that depend on continuity and differentiability are the same for both manifolds. These include the same Lyapunov exponents and fractal dimensions of the attractors on the manifold which are also connected by the manifolds’ diffeomorphism. Note, that many attractors are not themselves manifolds, but are contained in manifolds. However, we can talk about diffeomorphisms between attractors in the sense that TakensTakens (1981) has for time series.
The first fundamental property we consider is continuity. We would like to find a numerical/statistical test that could give us some idea if there is a continuous mapping between two attractor sampled data sets. We first developed a continuity statistic in previous papers Pecora, Carroll, and Heagy (1995); Pecora, Nichols, and Carroll (2007) and we introduce a more sophisticated version here using a null hypothesis to which we will give a geometric meaning.
@@
Once we choose a $\delta$ and gather $n_{\delta}$ points it holds we assume some test value of $\epsilon$. The $\epsilon$ is taken to be the radius of a sphere in $\cal N$. We now make the hypothesis that the $\delta$ points will fall at random into a sphere centered on $\phi({\bf x})$ that is twice the volume of the $\epsilon$-sized sphere. Then the probability $p_{\epsilon}$ that each of the $\delta$ points will fall in the $\epsilon$-sphere is 0.5. The probability that, say, $n_{\epsilon}$ of those points will fall into the $\epsilon$-sized sphere is given by the cumulative binomial distribution. If this latter probabilty is below 0.02, then we reject the null that the points fall into the double-volume set by chance. The natural steps here are to continue to shrink $\epsilon$ until the number of points are below a threshold where we cannot reject the null.
@@ VI Differentiability Statistic
For example, for chaotic or strange attractors we are not fully sampling the manifold containing the attractor. In fact some of those data points are sampled from a fractal object which cannot itself be a manifold. Hence, we are only examining a dynamically selected set of points of the manifold. But for the purposes of testing whether the RC is faithfully embedding the drive attractor, this should be enough.
@@
${\phi}^{-1}\circ{\phi}({\bf x}+{\bf h})=$ ${\phi}^{-1}\circ{\phi}({\bf x})+$
$D{\phi}^{-1}[D{\phi}\bm{\cdot}{\bf h}+\lVert{\bf h}\rVert\eta(\bf h)]$
$=$ ${\bf x}+D{\phi}^{-1}{\cdot}D{\phi}{\cdot}{\bf h}+$
@@
We see that to maintain the approximate equality we require $D{\phi}^{-1}{\cdot}D{\phi}=\bf 1$, where $\bf 1$ is the identity matrix. Thus, a diffeomorphism generates a tangent space mapping that is inverse to the tangent space mapping of the inverse diffeomorphim in each tangent space in the attractor’s manifold.
@@
Let $K$ be the dimension of the drive space and $NM$ be the dimension of the RC space. $N$ is the number of nodes and $M$ is the number of dynamical variables of each. Let $Q^{d}$ be the $L\times K$ matrix of $L$ local vectors around a fiducial point ${\bf x}_{0}$ in the drive. Let $Q^{r}$ be the $L\times NM$ matrix of local vectors around a fiducial point ${\bf r}_{0}$ in the RC. For reasons we explain below we choose $L$ to be an even number.
We first check that the number of singular values of $Q^{d}$ equals the number of singular values of $Q^{r}$. If not, then there cannot be a diffeomorphism as mentioned above. If $Q^{d}=Q^{r}$ we write $Q$ without superscripts for the local dimension of the drive and RC and we calculate the local linear maps between the drive and RC and check that they are indeed inverses of each other. To do this we pick at random 1/2 of the $L$ points and calculate the "forward" (drive $\rightarrow$ RC) $D\phi$ matrix. Then we use the remaining 1/2 of the $L$ points to calculate the "inverse" (RC $\rightarrow$ drive) $D\phi^{-1}$ matrix. We use different sets of points for each calculation to avoid the trivial result that the calculated $D\phi$ matrix would have an inverse since the points are merely "reversed", i.e. the domain and range of the points are exchanged. This explains why we choose $L$ to be at least twice the dimension of the drive. For better statistical sampling we also choose $L$ to be even larger.
@@
VII An important point about functional relations between two coupled attractors
@@
The solution to this situation is the continuity statistic itself. This is easy to see by assuming in Fig. 2 that we have a function $g:\mathcal{D}\rightarrow\mathcal{R}$. When we view $\mathcal{D}$ to be the domain of $g$ we search for small groups of points in $\mathcal{R}$ whose "partners" are in "local" sets of points. We will likely get points in $\mathcal{R}$ that are in in both branches of the period-2 attractor since we will be using the "Theiler window"Theiler (1986). This will lead to the detection of a "discontinuity" size that shows up in the continuity statistic approximately as the same $\epsilon^{*}$ at that attractor location no matter how many points we have in the data set. In this and any case where there are global "discontinuities" like this one will not see the trends we see later in this paper where there is evidence for a continuous function and $\epsilon^{*}$ does change in a predictable way as the number of data points in the time series changes.
VIII Attractor Comparison Statistic
@@
In this case we assume we have two time series that explore most of the attractors (labeled $A$ and $B$) for each, but are not sampled simultaneously. So the points in each are not time twins of each other, nor are they even of the same number (although the spatial dimensions must be the same). In this case we want a statistic that can distinguish whether two different time series are sampling the same or different attractors. To develop such a statistic we borrow concepts from set comparisons in metric spaces, which we modify to apply to data sampling of two or more sets rather than having access to all points in both sets, which would be impossible for sets containing a continuum of points.
@@
$d_{H}(A,B):=\text{max}\left\{\sup_{a\in A}\inf_{b\in B},\sup_{b\in B}\inf_{a\in A}\right\}$ (13)
This analytic form inspires our definition of set distances in data clouds, but while the form in Eq.(13) is necessary for mathematical proofs, the use of sup and inf is problematic for data analysis where a single point can distort the numerical analysis. For this reason we use a definition of the average Hausdorff dimension similar to usage in some data analysis (see, for example, Aydin et al. (2020).), but we keep the distances from A to B and B to A as separate statistics since one attractor can be arbitrarily small, but the other arbitrarily large. That is, the distances depend on the shape and placements of the sets with respect to each other and they are not necessarily even near equal. For example if A is small and placed near one side of B, but B is large and has points distant from A, then, on average, the distance from A to B will be small, but the distance from B to A will be large.
So, we replace inf with min and we use two distances to test whether A and B are close:
$d_{avg}(A,B):=\text{avg}_{a\in A}\>\min_{b\in B}\>d(a,B)$ (14)
$d_{avg}(B,A):=\text{avg}_{b\in B}\>\min_{a\in A}\>d(b,A),$
where we note that the order of $A$ and $B$ in the distance formula matters. We will see that these simple statistics are very helpful when we encounter multiple attractors in high-dimensional systems like an RC.
There is an important extension to the statistics in Eq.(14) that we can add in the case of time series from dynamical systems. We made the point earlier that for many time series the points nearby to each other in time will be well correlated. We can use this correlation (easily measured) to increase the power of Eq.(14) statistics. Points in time into the past or future of the closest points will also be close up to some fraction of the correlation time length. Hence, we can find these closest points by searching (in time) along the trajectories of attractors $A$ and $B$ add those points’ distances to the averages in Eq.(14), thus using the dynamics to make them a more reliable gauge of attractor identity. This is similar to Kennel’s use of "strands"Kennel and Abarbanel (2002). We do this in the following sections which use the attractor comparison statistic (ACS).
IX Numerical examples of diffeomorphic statistics

### v2

@@ I Introduction
What does all this have to do with reservoir computing systems? The main link there is provided by Takens’ embedding theorem Takens (1981) and work by Sauer, Yorke, and Casdagli Sauer, Yorke, and Casdagli (1991) and Hunt, Sauer, and Yorke Hunt and Yorke (1992). We will show a simplified explanation of how attractors on manifolds in dynamical systems driving a reservoir computer can (under the right circumstances) provide an embedding of the drive system in the reservoir system. The tests for this possibility using only data will involve the statistics we mentioned above. To get to this point we start by introducing reservoir systems and clarifying which principles are important for understanding their dynamics as well as which explanations of their behavior are misleading.
@@ II Reservoir Computing
Reservoir computing was introduced in the early 2000’s in a few papers Jaeger (2001); Maass, Natschlaeger, and Markram (2002); Jaeger (2007); H. Jaeger (2004), which influenced the machine learning field more than the nonlinear dynamics field, at least at first. To put it simply, it was initially shown that a reservoir computer (RC) could be driven by an input of one component of a dynamical system (the drive) and it could be trained to reproduce the other (unseen by the reservoir) drive components. And, once trained it could do this accurately for signals from other initial conditions of the same drive system.
@@ III Basic structure of Reservoir Computer and Test Systems
This process does not necessarily work for any input from the drive. Which ones are successful varies from drive to drive. The statistics we develop here can be useful in determining which combinations of drive signals work best. The outputs of the RC are signals from the nodes of the RC that are from the dynamical variables of the nodes. We extract all nodes’ signals, but it’s still somewhat an open question of which nodes of a subset might work just as well.
@@ IV Dynamics and Embeddings in a drive-
where ${\bf r}\in{\mathbbm{R}}^{NM}$. Using function composition iteratively in Eq.(2) and iterating on ${\bf r}(n)$ we can write Eq.(2) in an insightful way Jüngling . That is, we write,
@@
From the above Eq.(3) we see that the "next" time step ($n+1$) depends iteratively on the previous $n$ steps, assuming $n=0$ is the initial time and ${\bf{r}}(0)$ and ${\bf{x}}(0)$ the initial conditions. Eq.(3) is reminiscent of the form of Taken’s theorem Takens (1981) with respect to a dependence of the dynamics reconstruction on multiple time shifted terms.
Briefly, TakensTakens (1981) proved that for many dynamical systems one can reconstruct an entire attractor from a single measured time series quantity (even only one variable). This is done by using the measured variable (say, $x(t_{i})$ for many time points $\{t_{i}|i=1,2,...\}$) and constructing the $n$-dimensional vectors ${\bf v}(t_{i})=(x(t_{i}),x(t_{i-1}),x(t_{i-2}),...,x(t_{i-(n-1)}))$. The points ${\bf v}(t_{i})$ are then points on the attractor and this reconstruction is diffeomorphic to the actual attractor for non-chaotic systems. Since most chaotic attractors and some non-chaotic onesGrebogi et al. (1984) are fractal objects it is necessary to extend Takens proof using the concept of prevalenceHunt and Yorke (1992); Sauer, Yorke, and Casdagli (1991), which allows us to treat the Takens time-delay reconstruction as part of a manifold itself.
We can show that with a few more general assumptions Taken’s theory reconstruction indeed does apply here to RCs. In the case of continuous time evolution of the RC and drive we can get the same type of result as for Eq.(3), although it is a little more involved. Let $\psi_{t}$ be the flow of the drive system and $\Psi_{t}$ be the flow of the RC. The dynamics of the system is in the form of a skew product flow Kloeden and Yang (2021). That is the drive $\psi_{t}$ is autonomous, but the RC $\Psi_{t}$ is nonautonomous. The important property we will need is the cocycle property Kloeden and Yang (2021) of the flow on a nonautonomous system (the RC): $\forall\>s,t\in\mathbbm{R}^{+}\text{, }\bf{x}$ in the drive space, and $\bf{r}$ in the RC space, the following describes the flow:
@@
Then, iterating the above drive-RC embedding equations analogously to Eq.(3), Eq.(4) can be written as,
@@
The above explanation does not amount to a strict mathematical proof of the connection between the RC trajectory and some diffeomorphism of the drive attractor. We simply want to show that such a relation is reasonable and even likely. For more rigorous proofs of the RC-drive attractor relationship see the papers Allen Hart (2020); Grigoryeva, Hart, and Ortega (2021). Proofs for the theorems in these papers are at times limited to drive and RC systems with special properties (e.g. the necessity of contracting maps as a property of the dynamics). We emphasize that the above statements are not criticisms of the referenced papers. The proofs are involved and highly non-trivial. But so far the research on RCs suggests that the actual requirements for an embedding are less restrictive.
@@
In the next few sections we take the view that many drive-RC systems can generate time series data that can be used to test for the likelihood that the system has certain mathematical properties related to embeddings. This requires the development of statistics that we mentioned in the introduction.
@@ V Continuity Statistic
In this section on continuity and in the next section on differentiability, for simplicity of explanation we will treat the process of finding spatially nearby points as one of a simple search for the closest data points. This process is actually more subtle for dynamical time series. We will address the proper way to find nearby, but time-uncorrelated points.
Searching for closest points uses a metric on both spaces which we assume to be the usual square-root of sum of squares of component differences. This automatically translates into a set of open sets on each space as defined in the usual way by the metric and, hence, to a metric-induced topology on both drive and RC data spaces. As a result we use the same metric to define the properties of continuity, differentiability, etc. without resorting to a topological definition in terms of function inverses, open sets, etc. The latter would add a more abstract approach, but it would be unnecessarily so.
@@
Note that this type of mapping preserves a lot of geometric properties between the manifolds. Many properties of ${\mathcal{M}}$ are also true of ${\mathcal{N}}$ and vice versa. Especially important is the fact that many dynamical properties that depend on continuity and differentiability are the same for both manifolds. These include the same Lyapunov exponents and fractal dimensions of the attractors on the manifold which are also connected by the manifolds’ diffeomorphism.
Note, that many attractors are not themselves manifolds, but are contained in manifolds. As mentioned in the previous section, the prevalence extension to Takens theorem allows us to estimate local properties like continuity and differentiability, for which the latter needs a tangent space on the manifold.
The first fundamental property we consider is continuity. We would like to find a numerical/statistical test that could give us some idea if there is a continuous mapping between two simultaneously sampled attractor data sets. We first developed a continuity statistic in previous papers Pecora, Carroll, and Heagy (1995); Pecora, Nichols, and Carroll (2007) and we introduce a more sophisticated version here using a null hypothesis to which we will give a geometric meaning.
@@
Once we choose a $\delta$ and gather the $n_{\delta}$ points it holds we assume some test value of $\epsilon$. The $\epsilon$ is the radius of a sphere in $\cal N$. We now make the hypothesis that the $\delta$ points will fall at random into a sphere centered on $\phi({\bf x})$ that is twice the volume of the $\epsilon$-sized sphere. Then the probability $p_{\epsilon}$ that each of the $\delta$ points will fall in the $\epsilon$-sphere is 0.5. The probability that, say, $n_{\epsilon}$ of those points will fall into the $\epsilon$-sized sphere is given by the cumulative binomial distribution. If this latter probabilty is below 0.02, then we reject the null that the points fall into the double-volume set by chance. The natural steps here are to continue to shrink $\epsilon$ and $\delta$, as necessary, until the number of points are below a threshold where we cannot reject the null.
@@ VI Differentiability Statistic
${\phi}^{-1}\circ{\phi}({\bf x}+{\bf h})=$ ${\bf x}+D{\phi}^{-1}{\cdot}D{\phi}{\cdot}{\bf h}+$
@@
We see that to maintain the approximate equality for all small $\bf{h}$ we require $D{\phi}^{-1}{\cdot}D{\phi}=\bf 1$, where $\bf 1$ is the identity matrix. Thus, a diffeomorphism generates a tangent space mapping that is inverse to the tangent space mapping of the inverse diffeomorphim in each tangent space in the attractor’s manifold.
@@
Let $K$ be the dimension of the drive space and $NM$ be the dimension of the RC space. $N$ is the number of nodes and $M$ is the number of dynamical variables of each. Let $Q^{d}$ be the $L\times K$ matrix of $L$ local vectors around a fiducial point ${\bf x}_{0}$ in the drive. Let $Q^{r}$ be the $L\times NM$ matrix of local vectors around a fiducial point ${\bf r}_{0}$ in the RC. For reasons we explain below we choose $L$ to be an even number $\geq$ the number of singular values of $Q^{r}$ and $Q^{d}$.
We first check that the number of singular values $P^{d}$ of $Q^{d}$ equals the number of singular values $P^{r}$ of $Q^{r}$. If not, then there cannot be a diffeomorphism as mentioned above. If $P^{d}=P^{r}$ we write $P$ without superscripts for the local dimension of the drive and RC and we calculate the local linear maps between the drive and RC and check that they are indeed inverses of each other. To do this we pick at random 1/2 of the $L$ points and calculate the "forward" (drive $\rightarrow$ RC) $D\phi$ matrix. Then we use the remaining 1/2 of the $L$ points to calculate the "inverse" (RC $\rightarrow$ drive) $D\phi^{-1}$ matrix. We use different sets of points for each calculation to avoid the trivial result that the calculated $D(\phi)$ matrix would have an inverse, $D(\phi^{-1})$ since the points are merely "reversed", i.e. the domain and range of the points are exchanged. This explains why we choose $L$ to be at least twice $P$, the dimension of the drive. For better statistical sampling we also choose $L$ to be even larger.
@@
What remains is augmenting the above analysis with a null hypothesis as we did with the continuity statistic. This means asking if the result of the product of the linear tangent maps being near to $\bf 1$ could happen "accidentally." At this point we have not yet developed such a null test, so we will rely on the statistic being much less than 1.0 as evidence of a diffeomorphism.
VII One directional differentiability
Here we show a statistic for testing functional differentiability in one direction only. We will not use this in the reservoir examples in this paper, but we show an example of it in the supplemental material about the statistics of continuous, but not differentiable functions.
To develop this statistic we borrow from the diffeomorphism statistic. We assume that we have a function $\phi$ which is differentiable in only one direction (from domain to range) as defined in Eq.(10). We use SVD to find the dimension $P$ of the tangent space at a fiducial point $\bf x$ and the dimension of the tangent space in the range at the image of $\bf x$. As before, we find an even number of points ($L>P$) at a fiducial point $\bf x$. We use $L/2$ points to find a local matrix approximation for $D\phi$.
The system of points from the domain (the drive here) to the range (the RC here) has to be equal- or over-determinedGolub and Van Loan (2013). This is not to just deal with possible noise in the measured time series (since we are considering that the time series may come from measurements), but also that we cannot have a system that is underdetermined, but has a local linear map from the domain to the range. That is, the dimension of the manifold in the domain must be less than or equal to the dimension in the range. The SVDs of the local domain and range manifolds test for this and for determining the form (dimensionality) of the local derivative.
We get the approximate $D\phi$ using the usual least squares approach Golub and Van Loan (2013) to map the remaining points to the range of the function and compare these vectors to the actual vectors in the range using root mean square average differences, expressed as a fraction of the actual vector magnitudes. Small values of these fractions suggest that the function from domain to range is a differentiable function. We judge the function’s differentiability on the domain by averaging over the fractional errors in the vectors. As with the differentiability statistic we would eventually like to develop a null hypothesis to aid in understanding the meaning of the statistic better.
VIII An important point about functional relations between two coupled attractors
@@
The solution to this situation is the continuity statistic itself. This is easy to see by assuming in Fig. 2 that we have a function $g:\mathcal{D}\rightarrow\mathcal{R}$. When we view $\mathcal{D}$ to be the domain of $g$ we search for small groups of points in $\mathcal{R}$ whose "partners" are in "local" sets of points. We will likely get points in $\mathcal{R}$ that are in in both branches of the period-2 attractor since we will be using the "Theiler window"Theiler (1986). This will lead to the detection of a "discontinuity" size that shows up in the continuity statistic approximately as the same $\epsilon^{*}$ at that attractor location no matter how many points we have in the data set. In this and any case where there are global "discontinuities" like this one we will not see the trends we see later in this paper where there is evidence for a continuous function and $\epsilon^{*}$ does change in a predictable way as the number of data points in the time series changes.
IX Attractor Comparison Statistic
@@
In this case we assume we have two time series that explore most of the attractors (labeled $\mathcal{A}$ and $\mathcal{B}$) for each, but are not sampled simultaneously. So the points in each are not time twins of each other, nor are they even of the same number (although the spatial dimensions must be the same). In this case we want a statistic that can distinguish whether two different time series are sampling the same or different attractors. To develop such a statistic we borrow concepts from set comparisons in metric spaces, which we modify to apply to data sampling of two or more sets rather than having access to all points in both sets, which would be impossible for sets containing a continuum of points.
@@
$d_{H}(\mathcal{A},\mathcal{B}):=\text{max}\left\{\sup_{a\in\mathcal{A}}\inf_{b\in\mathcal{B}},\sup_{b\in\mathcal{B}}\inf_{a\in\mathcal{A}}\right\}$ (13)
This analytic form inspires our definition of set distances in data clouds, but while the form in Eq.(13) is necessary for mathematical proofs, the use of sup and inf is problematic for data analysis where a single point can distort the numerical analysis. For this reason we use a definition of the average Hausdorff dimension similar to usage in some data analysis (see, for example, Aydin et al. (2020).), but we keep the distances from $\mathcal{A}$ to $\mathcal{B}$ and $\mathcal{B}$ to $\mathcal{A}$ as separate statistics since one attractor can be arbitrarily small, but the other arbitrarily large. That is, the distances depend on the shape and placements of the sets with respect to each other and they are not necessarily even near equal. For example if $\mathcal{A}$ is small and placed near one side of $\mathcal{B}$, but $\mathcal{B}$ is large and has points distant from $\mathcal{A}$, then, on average, the distance from $\mathcal{A}$ to $\mathcal{B}$ will be small, but the distance from $\mathcal{B}$ to $\mathcal{A}$ will be large.
So, we replace inf with min and we use two distances to test whether $\mathcal{A}$ and $\mathcal{B}$ are close:
$d_{avg}(\mathcal{A},\mathcal{B}):=\text{avg}_{a\in\mathcal{A}}\>\min_{b\in\mathcal{B}}\>d(a,\mathcal{B})$ (14)
$d_{avg}(\mathcal{B},\mathcal{A}):=\text{avg}_{b\in\mathcal{B}}\>\min_{a\in\mathcal{A}}\>d(b,\mathcal{A}),$
where the order of $\mathcal{A}$ and $\mathcal{B}$ in the distance formula matters. We will see that these simple statistics are very helpful when we encounter multiple attractors in high-dimensional systems like an RC.
There is an important extension to the statistics in Eq.(14) that we can add in the case of time series from dynamical systems. We made the point earlier that for many time series the points nearby to each other in time will be well correlated. We can use this correlation (easily measured) to increase the power of Eq.(14) statistics. Points in time into the past or future of the closest points will also be close up to some fraction of the correlation time length. Hence, we can find these closest points by searching (in time) along the trajectories of attractors $\mathcal{A}$ and $\mathcal{B}$ add those points’ distances to the averages in Eq.(14), thus using the dynamics to make them a more reliable gauge of attractor identity. This is similar to Kennel’s use of "strands"Kennel and Abarbanel (2002). We do this in the following sections which use the attractor comparison statistic (ACS).
X Numerical examples of diffeomorphic statistics